William throws the ball to Caleb who is standing 40 meters away. 1/4 of the way the ball is going Caleb it’s height is at 8 meters. Calculate the max height.
1
Expert's answer
2019-05-15T10:12:07-0400

Solution. Consider the general equations of motion at an angle where the movement in the horizontal and vertical direction is independent. In the horizontal direction, the body moves at a constant speed.


"v_x=vcos\alpha"

where v is initial velocity of the ball.

In the vertical direction, the body moves with an initial upward curvature and the equation of skrasty and height can be written as


"v_y=vsin\alpha - gt"

"h=vtsin\alpha -\frac {gt^2} {2}"

where g=9.8 m/s^2 is constant gravity; t is time. According to the condition of the problem


"vt_1cos\alpha=10"

"8=vt_1sin\alpha -\frac {gt_1^2} {2}"

Therefore


"t_1=\frac {10} {vcos\alpha}"

"8=10tan\alpha -\frac {100g} {2v^2cos^2\alpha}"

The ball will be at the maximum height at exactly half of its trajectory. Hence time is equal to


"t_2=\frac {20} {vcos \alpha}."

On the other hand, the maximum lift height corresponds to zero vertical speed.


"0=vsin\alpha - gt_2"

"t_2=\frac {vsin\alpha} {g}"

Therefore


"v^2 sin \alpha cos \alpha =20g"

As result

"8=10tan\alpha -\frac {5v^2 sin \alpha cos \alpha} {2v^2cos^2\alpha}"

"8=10tan\alpha-\frac {5} {2} tan\alpha \Longrightarrow tan\alpha=\frac {16} {15}"

The maximum lift height will be found using the formula


"h_{max}=\frac {v^2sin^2 \alpha} {2g}"

Hence


"h_{max}=\frac {v^2sin^2 \alpha} {\frac {v^2 sin \alpha cos \alpha} {10}}=\frac {10v^2sin^2 \alpha} {v^2 sin \alpha cos \alpha}=10tan\alpha"

"h_{max}=10 \times \frac {16} {15}=\frac {32} {3} meters"



Answer.

"h_{max}=\frac {32} {3} meters"